OpenStudy (anonymous):

Sebastian's parents agree to lend him 200 for some gaming equipment, as long as he pays them back 250 in a year's time. What is the rate of simple annual interest in this arrangement?

OpenStudy (anonymous):

Simple interest = P*I*N where P= Original amount borrowed, I = interest rate and N = the amount of years

OpenStudy (anonymous):

K then. The interest is 50 dollars, cause that's what's he paying extra, and $50 is 25% of 200. That's what I believe it is.

OpenStudy (anonymous):

\[50=200*I*1\]

OpenStudy (anonymous):

yeah i did compound interest... i'm not really into the economic math

OpenStudy (anonymous):

so its 50%

OpenStudy (anonymous):

I=50/200

OpenStudy (anonymous):

No. @trisha30 50/200=1/4 so it's 25%

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

No prob.